# warped area minimal hypersurface and warped product metric

### 13 Mar 2026 v1

@@ 1. Introduction
A classic rigidity result concerning area-minimizing hypersurfaces, due to Cai-Galloway[CG00], asserts the following: if a three-dimensional manifold $(M,g)$ has non-negative scalar curvature ($\operatorname{Sc}_{g}\geq 0$) and contains a two-sided, locally area-minimizing torus $\Sigma$, then $M$ is flat in a neighborhood of $\Sigma$; equivalently, $M$ admits a product metric near $\Sigma$. This confirms a conjecture originally proposed by Fischer-Colbrie-Schoen[FCS80]. Building on the results of Schoen-Yau[SY79] on the classifications of area-minimizing surfaces in three-dimensional manifolds with positive scalar curvature, any such surface in $M$ must be homeomorphic to either $\mathbb{S}^{2}$ (the 2-sphere) or $\mathbb{RP}^{2}$ (the real projective plane). Subsequently, Bray-Brendle-Neves[BBN10] investigated the case of an area-minimizing 2-sphere $\Sigma$ in a compact three-dimensional manifold $M^{3}$ with scalar curvature bounded below by 2. In particular, they proved that the area of $\Sigma$ is at most $4\pi$; crucially, equality holds if and only if $M^{3}$ is covered by $\mathbb{S}^{2}\times\mathbb{R}$. This implies that, in the rigid case, $M$ locally admits a product metric near $\Sigma$. Since then, additional rigidity results establishing local product metric splittings for area-minimizing hypersurfaces have been obtained [CEM19],[Zhu20],[CLZ25].
A natural question thus arises: under what conditions do hypersurfaces admit local splittings as warped product metrics? In this paper, we address this question and establish several key results. We further elucidate the geometric significance of warped(or weighted) area-minimizing hypersurfaces. To begin with, we analyze the Ricci curvature and scalar curvature of the warped product manifold $(\mathbb{S}^{1}\times N^{n-1},g=dt^{2}+f^{2}(t)g_{N})$ where $f$ denotes a positive smooth function on $\mathbb{S}^{1}$ and $(N^{n-1},g_{N})$ is a Riemannian manifold. The Ricci curvature of the metric $g$ is
@@
Therefore, we study the warped minimal hypersurface under the spectral Ricci curvature condition and the spectral scalar curvature condition.
@@
Throughout this paper, we let $\gamma=n-1$.
@@
with $a<\frac{4}{n-1}$. Catino-Mari-Mastrolia-Roncoroni[CMMR25] and Antonelli-Pozzetta-Xu[APX24] proved that $M$ splits as a product manifold. As Antonelli-Pozzetta-Xu pointed out, such a product metric splitting fails to hold when $a\geq\frac{4}{n-1}$. Nevertheless, for the case $n=3$, if there exists a compact warped area-minimizing hypersurface, then the manifold $M^{3}$ with two ends will locally split as a warped product metric according to Theorem 1.2. For nonnegative Bakry–Émery Ricci curvature, Liu [Liu13] obtains a locally split product metric by studying the weighted area-minimizing hypersurface.
@@
where $u$ is a smooth positive function on $M$ and $5\leq n\leq 7$. If $H_{n-1}(M,\mathbb{Z})\neq 0$ and $0\neq[\Sigma]\in H_{n-1}(M,\mathbb{Z})$ such that there is no positive scalar curvature on $\Sigma$, then, $u$ is constant on $\Sigma$ and $(M^{n},g)$ locally splits as $(M^{n},g=dt^{2}+u^{-2}g_{\Sigma})$. Moreover, $(\Sigma,g_{\Sigma})$ is Ricci flat.
@@
The operator $\gamma\Delta_{g}+\frac{1}{2}\operatorname{Sc}_{g}$ is greater than the conformal operator $\frac{2(n-1)}{n-2}\Delta_{g}+\frac{1}{2}\operatorname{Sc}_{g}$ for $5\leq n\leq 7$. That is, if the $\gamma\Delta_{g}+\frac{1}{2}\operatorname{Sc}_{g}$ is positive, we can not use the conformal change to obtain a positive scalar curvature metric.
@@ 2. Basic fact about warped or weighted minimal hypersurface
where $\gamma=n-1$ and $\mathcal{H}^{n-1}$ denotes the (n1)(n-1) dimension Hausdorff measure induced by the metric $g$. Let $H$ and $A$ stand for the mean curvature and the second fundamental of $\Sigma$, respectively.
Lemma 2.1.
@@
We say $\Sigma$ a warped(or weighted) area-minimizing hypersurface if $H+\gamma u^{-1}u_{\nu}=0$ along $\Sigma$.
@@
Consider a warped area-minimizing hypersurface $\Sigma$. Assume that $\Sigma_{t}$ is a smooth $1$-parameter family of smooth hypersurfaces along the variation vector field $Y$ and normal speed $\phi=\langle Y,\nu\rangle$ at $t=0$. Then
@@
A warped area-minimizing hypersurface $\Sigma$ is called stable if
@@ 4. Spectral scalar curvature case
The first step is to prove that $H_{t}+\gamma\omega_{\nu_{t}}=0$ on $\Sigma_{t}$ for $t\in(-\epsilon,\epsilon)$. Then the stable warped area-minimizing property of $\Sigma_{t}$ follows. We do that in Proposition 4.3 and $4.$ ∎
@@
Using the spectral scalar curvature bound, and with suitable grouping of terms, we see
@@
$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}$
@@
$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla_{\Sigma}u|^{2}$
$\quad-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}+\frac{n}{2(n-1)}\tilde{H}^{2}(t)+\tilde{H}(t)\left(\gamma w_{\nu_{t}}-\frac{n\gamma}{n-1}w_{\nu_{t}}\right)$
@@
$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla_{\Sigma}u|^{2}$
$-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}+\tilde{H}(t)\left(-w_{\nu_{t}}\right).$
@@
$\geq$ $\phi_{t}^{-1}\tilde{H}^{\prime}(t)-\tilde{H}(t)w_{\nu_{t}}-\tfrac{1}{2}\operatorname{Sc}_{\Sigma}.$
@@
$=$ $\tfrac{\gamma(\gamma+2)}{4}u^{-2}|\nabla_{\Sigma}u|^{2}-\tfrac{\gamma}{2}u^{-1}\Delta_{\Sigma_{t}}u-\gamma\langle\nabla_{\Sigma}u,\nabla_{\Sigma_{t}}\xi_{t}\rangle u^{-1}$
@@
$=$ $\tfrac{\gamma(2-\gamma)}{4}u^{-2}|\nabla_{\Sigma}u|^{2}+\tfrac{\gamma}{2}u^{-1}\Delta_{\Sigma_{t}}u+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}$
$=$ $\tfrac{\gamma(4-\gamma)}{4}|\nabla_{\Sigma}w|^{2}+\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}.$
@@
$=$ $-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}.$
@@
$\tilde{H}^{\prime}(t)-\tilde{H}(t)w_{\nu_{t}}\phi_{t}\leq\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\right)\phi_{t}.$
@@
$\leq$ $\frac{\int_{\Sigma_{t}}\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\right)\phi_{t}\varphi_{t}}{\int_{\Sigma_{t}}\varphi_{t}}$
@@
$\int_{\Sigma_{t}}\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\right)\phi_{t}\varphi_{t}>0.$
@@
$\int_{\Sigma_{t}}\left(|\nabla_{\Sigma}\psi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\psi_{t}^{2}\right)+\int_{\Sigma}\gamma\psi_{t}\langle\nabla_{\Sigma_{t}}w,\nabla_{\Sigma_{t}}\psi_{t}\rangle+\int_{\Sigma_{t}}\tfrac{\gamma(2-\gamma)}{4}|\nabla_{\Sigma}w|^{2}>0.$
@@
(4.5) $\int_{\Sigma}\left[\frac{2(n-2)}{n-3}|\nabla_{\Sigma}\psi_{t}|^{2}+\frac{1}{2}\operatorname{Sc}_{\Sigma}\psi_{t}^{2}\right]>0,$
@@
In the following, we deduce that the metric $g$ can be locally written as a warped product metric. From lemma 4.1 and lemma 4.2, we can obtain that, locally, $M^{n}$ can be written as $(-\epsilon,\epsilon)\times\Sigma$ such that each $\Sigma_{t}$ is a warped area-minimizing hypersurface and $\Sigma_{t}$ is umbilical. Then the metric $g$ can be written as $g=dt^{2}+f^{2}(t)g_{\Sigma}$ for some positive function $f$ on $(-\epsilon,\epsilon)$. Since $H_{t}=-(n-1)u^{-1}u^{\prime}(t)=(n-1)f^{-1}f^{\prime}(t)$, we obtain $f(t)=u^{-1}(t)$. By an open and closed argument, by passing to a suitable cover $\tilde{M}$, $(\tilde{M}^{n},g)$ is isometric to $(\mathbb{S}^{1}\times\Sigma^{n-1},ds^{2}+f^{2}(t)g|_{\Sigma})$. By the equation (1.3), we can obtain that $(\Sigma,g_{\Sigma})$ is scalar flat. Because $\Sigma$ admits no positive scalar curvature, we can deduce that $(\Sigma,g_{\Sigma})$ is Ricci flat.

### 21 Mar 2026 v2

# warped area-minimizing hypersurface and warped product metric

@@ 1. Introduction
For compact manifolds, a classic rigidity result concerning area-minimizing hypersurfaces, due to Cai-Galloway[CG00], asserts the following: if a three-dimensional manifold $(M,g)$ has non-negative scalar curvature ($\operatorname{Sc}_{g}\geq 0$) and contains a two-sided, locally area-minimizing torus $\Sigma$, then $M$ is flat in a neighborhood of $\Sigma$; equivalently, $M$ admits a product metric near $\Sigma$. This confirms a conjecture originally proposed by Fischer-Colbrie-Schoen[FCS80]. Building on the results of Schoen-Yau[SY79] on the classifications of area-minimizing surfaces in three-dimensional manifolds with positive scalar curvature, any such surface in $M$ must be homeomorphic to either $\mathbb{S}^{2}$ (the 2-sphere) or $\mathbb{RP}^{2}$ (the real projective plane). Subsequently, Bray-Brendle-Neves[BBN10] investigated the case of an area-minimizing 2-sphere $\Sigma$ in a compact three-dimensional manifold $M^{3}$ with scalar curvature bounded below by 2. In particular, they proved that the area of $\Sigma$ is at most $4\pi$; crucially, equality holds if and only if $M^{3}$ is covered by $\mathbb{S}^{2}\times\mathbb{R}$. This implies that, in the rigid case, $M$ locally admits a product metric near $\Sigma$. Later, Zhu[Zhu20] extended Bray-Brendle-Neves’ result to a more general setting for an $n$-dimensional manifold $M^{n}$ that admits a non-zero degree map to $\mathbb{S}^{2}\times\mathbb{T}^{n}$, where $3\leq n+2\leq 7$. Recently, in the rigid case of their theorem, Chu-Lee-Zhu[CLZ25] proved that the universal cover of an $n$-dimensional manifold $(M^{n},g)$ with non-trivial homology group $H_{n}(M)$ splits isometrically as a product, assuming the bi-Ricci curvature is bounded from below by n1n-1.
For noncompact manifolds, there is also an extensive literature concerning metric splitting theorems in the presence of area-minimizing hypersurfaces. For instance, we refer the reader to [AR89],[LW04],[CEM19],[CCE16],[HSY25],[Zhu23].
A natural question thus arises: under what conditions do hypersurfaces admit local splitting as warped product metrics? In this paper, we address this question for compact manifolds and establish several key results. We further elucidate the geometric significance of warped(or weighted) area-minimizing hypersurfaces. To begin with, we analyze the Ricci curvature and scalar curvature of the warped product manifold $(\mathbb{S}^{1}\times N^{n-1},g=dt^{2}+f^{2}(t)g_{N})$ where $f$ denotes a positive smooth function on $\mathbb{S}^{1}$ and $(N^{n-1},g_{N})$ is a Riemannian manifold. The Ricci curvature of the metric $g$ is
@@
The equations (1.4) and (1.5) are similar to the spectral curvature. Therefore, we study the warped minimal hypersurface under the spectral Ricci curvature condition and the spectral scalar curvature condition.
@@
Spectral curvature plays an important role in the study of the stable Bernstein problem; see [CLMS24],[Maz24]. A variety of geometric results have been established concerning spectral curvature, including width estimates [HKKZ24],[CS25], diameter and volume estimates [CR21], [AX24],[CH26],[Wu26] and splitting theorems(product case) [APX24],[CMMR25],[HW26],[Yeu25].
We obtain some splitting theorems(warped product case) under the spectral curvature condition as follows. Throughout this paper, we let $\gamma=n-1$.
@@
with $a<\frac{4}{n-1}$. Catino-Mari-Mastrolia-Roncoroni[CMMR25] and Antonelli-Pozzetta-Xu[APX24] proved that $M$ splits as a product manifold. As Antonelli-Pozzetta-Xu pointed out, such a product metric splitting fails to hold when $a\geq\frac{4}{n-1}$. Nevertheless, for the case $n=3$ and a=2, if there exists a compact warped area-minimizing hypersurface, then the manifold $M^{3}$ with two ends will locally split as a warped product metric according to Theorem 1.2. For nonnegative Bakry–Émery Ricci curvature, Liu [Liu13] obtains a locally split product metric by studying the weighted area-minimizing hypersurface. Li-Wang[LW06] also proved that the manifold splits as a warped product under a lower Ricci curvature bound and additional geometric conditions.
@@
where $u$ is a smooth positive function on $M$ and $4\leq n\leq 7$. If $H_{n-1}(M,\mathbb{Z})\neq 0$ and $0\neq[\Sigma]\in H_{n-1}(M,\mathbb{Z})$ such that there is no positive scalar curvature on $\Sigma$, then, $u$ is constant on $\Sigma$ and $(M^{n},g)$ locally splits as $(M^{n},g=dt^{2}+u^{-2}g_{\Sigma})$. Moreover, $(\Sigma,g_{\Sigma})$ is Ricci flat.
@@
If we let $v=u^{\frac{n-2}{2}}$, then the inequality (1.10) implies that
$\frac{2(n-1)}{n-2}v^{-1}\Delta_{g}v+\frac{1}{2}\operatorname{Sc}_{g}\geq 0.$
Then we can use the conformal change to obtain a nonnegative scalar curvature metric.
@@
Acknowledgements: The author would like to express his gratitude to Professor Thomas Richard and Professor Hong Han for their interest and conversation about this work.
@@ 2. Basic fact about warped or weighted minimal hypersurface
where $\gamma=n-1$ and $\mathcal{H}^{n-1}$ denotes the (n1)(n-1) dimension Hausdorff measure induced by the metric $g$. Let $H$ and AA stand for the mean curvature and the second fundamental of $\Sigma$, respectively. First, we give the definition of the warped area-minimizing hypersurface.
Definition 2.1.
Let $\Sigma$ be a compact hypersurface in a Riemannian manifold $(M,g)$. Suppose that the homology class $\Sigma]\in H_{n-1}(M,\mathbb{Z})$ of $\Sigma$ is non-zero, i.e., $\Sigma]\neq 0$ in $H_{n-1}(M,\mathbb{Z})$. Then $\Sigma$ is called a warped area-minimizing hypersurface if
$\int_{\Sigma}u^{\gamma}\mathrm{d}\mathcal{H}^{n-1}=\inf_{\hat{\Sigma}\in[\Sigma]}\{\int_{\hat{\Sigma}}u^{\gamma}\mathrm{d}\mathcal{H}^{n-1}\}.$
Lemma $2.$
@@
We say $\Sigma$ a warped(or weighted) minimal hypersurface if $H+\gamma u^{-1}u_{\nu}=0$ along $\Sigma$.
@@
Consider a warped minimal hypersurface $\Sigma$. Assume that $\Sigma_{t}$ is a smooth 11-parameter family of smooth hypersurfaces along the variation vector field $Y$ and normal speed $\phi=\langle Y,\nu\rangle$ at $t=0$. Then
@@
A warped minimal hypersurface $\Sigma$ is called stable if
@@ 4. Spectral scalar curvature case
The first step is to prove that $H_{t}+\gamma\omega_{\nu_{t}}=0$ on $\Sigma_{t}$ for $t\in(-\epsilon,\epsilon)$. Then the stable warped area-minimizing property of $\Sigma_{t}$ follows. We do that in Proposition 4.3 and 4.4. ∎
@@
Using the spectral scalar curvature bound and (4.4) with suitable grouping of terms, we see
@@
$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})-\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}$
@@
$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla_{\Sigma_{t}}u|^{2}$
$\quad-\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}+\frac{n}{2(n-1)}\tilde{H}^{2}(t)+\tilde{H}(t)\left(\gamma w_{\nu_{t}}-\frac{n\gamma}{n-1}w_{\nu_{t}}\right)$
@@
$\geq$ $(\phi_{t}^{-1}\Delta_{\Sigma_{t}}\phi_{t}+\gamma u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\phi_{t}^{-1})+\frac{\gamma(\gamma-3)}{2}u^{-2}|\nabla_{\Sigma_{t}}u|^{2}$
$-\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}+\tilde{H}(t)\left(-w_{\nu_{t}}\right).$
@@
$\geq$ $\phi_{t}^{-1}\tilde{H}^{\prime}(t)-\tilde{H}(t)w_{\nu_{t}}-\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}.$
@@
$=$ $\tfrac{\gamma(\gamma+2)}{4}u^{-2}|\nabla_{\Sigma_{t}}u|^{2}-\tfrac{\gamma}{2}u^{-1}\Delta_{\Sigma_{t}}u-\gamma\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\xi_{t}\rangle u^{-1}$
@@
$=$ $\tfrac{\gamma(2-\gamma)}{4}u^{-2}|\nabla_{\Sigma_{t}}u|^{2}+\tfrac{\gamma}{2}u^{-1}\Delta_{\Sigma_{t}}u+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}$
$=$ $\tfrac{\gamma(4-\gamma)}{4}|\nabla_{\Sigma_{t}}w|^{2}+\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w+\Delta_{\Sigma_{t}}\xi_{t}+|\nabla_{\Sigma_{t}}\xi_{t}|^{2}.$
@@
$-\phi_{t}^{-1}\left(\Delta_{\Sigma_{t}}\phi_{t}+\gamma\phi_{t}u^{-1}\Delta_{\Sigma_{t}}u+\gamma u^{-1}\langle\nabla_{\Sigma_{t}}u,\nabla_{\Sigma_{t}}\phi_{t}\rangle\right)-\frac{\gamma(\gamma-3)}{2}|\nabla_{\Sigma_{t}}w|^{2}$
$=$ $-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma_{t}}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}.$
@@
$\tilde{H}^{\prime}(t)-\tilde{H}(t)w_{\nu_{t}}\phi_{t}\leq\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma_{t}}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}\right)\phi_{t}.$
@@
$\leq$ $\frac{\int_{\Sigma_{t}}\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma_{t}}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}\right)\phi_{t}\varphi_{t}}{\int_{\Sigma_{t}}\varphi_{t}}$
@@
$\int_{\Sigma_{t}}\left(-\tfrac{\gamma(\gamma-2)}{4}|\nabla_{\Sigma_{t}}w|^{2}-\tfrac{\gamma}{2}\Delta_{\Sigma_{t}}w-\Delta_{\Sigma_{t}}\xi_{t}-|\nabla_{\Sigma_{t}}\xi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma_{t}}\right)\phi_{t}\varphi_{t}>0.$
@@
$\int_{\Sigma_{t}}\left(|\nabla_{\Sigma_{t}}\psi_{t}|^{2}+\tfrac{1}{2}\operatorname{Sc}_{\Sigma}\psi_{t}^{2}\right)+\int_{\Sigma}\gamma\psi_{t}\langle\nabla_{\Sigma_{t}}w,\nabla_{\Sigma_{t}}\psi_{t}\rangle+\int_{\Sigma_{t}}\tfrac{\gamma(2-\gamma)}{4}|\nabla_{\Sigma_{t}}w|^{2}>0.$
@@
(4.5) $\int_{\Sigma}\left[\frac{2(n-2)}{n-3}|\nabla_{\Sigma_{t}}\psi_{t}|^{2}+\frac{1}{2}\operatorname{Sc}_{\Sigma_{t}}\psi_{t}^{2}\right]>0,$
@@
In the following, we deduce that the metric $g$ can be locally written as a warped product metric. From lemma 4.1 and lemma 4.2, we can obtain that, locally, $M^{n}$ can be written as $(-\epsilon,\epsilon)\times\Sigma$ such that each $\Sigma_{t}$ is a warped area-minimizing hypersurface and $\Sigma_{t}$ is umbilical. Then the metric $g$ can be written as $g=dt^{2}+f^{2}(t)g_{\Sigma}$ for some positive function $f$ on $(-\epsilon,\epsilon)$. Since $H_{t}=-(n-1)u^{-1}u^{\prime}(t)=(n-1)f^{-1}f^{\prime}(t)$, we obtain $f(t)=u^{-1}(t)$. By an open and closed argument, by passing to a suitable cover $\tilde{M}$, $(\tilde{M}^{n},g)$ is isometric to $(\mathbb{S}^{1}\times\Sigma^{n-1},ds^{2}+f^{2}(t)g|_{\Sigma})$. Because $(\Sigma,g_{\Sigma})$ is scalar flat and $\Sigma$ admits no positive scalar curvature, we can deduce that $(\Sigma,g_{\Sigma})$ is Ricci flat.